A capacitor with double the area and double the dielectric thickness will have.

Option A. double the capacitance.
Option B. the same capacitance.
Option C. half the capacitance.
1
Expert's answer
2019-10-03T09:39:00-0400


The original capacitance is given by


"C_{o}=k\frac{\varepsilon_{o}A_{o}}{d_{o}}"


Where

  • Dielectric constant "K"
  • Parallel plate area "A_{o}"
  • distance between more plates "d_{o}"



Considering two dielectrics of constanta K (equal) and width of "do"


Each half can be considered a capacitor, if it is assumed that the plates have an electric charge + Q and -Q, the electric field in each capacitor is:


"E_{1}=\frac{Q}{k\varepsilon_{o} *2A_{o}} \\ E_{2}=\frac{Q}{k\varepsilon_{o} *2A_{o}}" remember that you have twice the area Ao


The potential difference between the plates is equal to:


remember that each capacitor has a thickness do

"V=E_{1}d_{o}+E_{2}d_{o} \\ V=\frac{Q}{k\varepsilon_{o} *2A_{o}}+\frac{Q}{k\varepsilon_{o} *2A_{o}} \\ V=2\frac{Q}{k\varepsilon_{o} *2A_{o}} \\V=\frac{Q}{k\varepsilon_{o} *A_{o}}"


The new capacitance is equal to


"C=\frac{Q}{V} \\ C=\frac{Q}{\frac{Q}{k\varepsilon_{o} *A_{o}}}"


Simplifying the new capacitance is"C=k\varepsilon_{o}A_{o}"



Comparing with the original capacitance

"\frac{C}{C_{o}}=\frac{k\varepsilon_{o}A_{o}}{k\frac{\varepsilon_{o}A_{o}}{d_{o}}}\\\frac{C}{C_{o}}=1 \\C=C_{o}"


Solution:"\boxed{\text{Option B. the same capacitance.}}"